### 18 Dec 2023 v1

@@ 5.5 Explainable patient activity recognition based on interpretable models
One of the major hindrances in generating spatio-temporal interpretation outputs is the lack of a model-agnostic interpretation pipeline. It is challenging, as the spatio-temporal interpretation illustrates the significance of detecting biases in video data and action associations through deeper network analysis. The motion-guided sampler (MGSampler) [242] attempted to interpret the dynamic motion characteristics by leveraging motion as a fundamental signal for frame selection. It utilizes a frame selection approach that ensures an even distribution of frames across important segments, characterized by high motion salience, achieved through the accumulation of motion data. M-MiT [143] provides a more comprehensive method that is focused on interpreting action recognition models. This method is focused on examining these interpretable features and how they contribute to a network’s output. It reveals insights that might otherwise remain concealed without the capacity to identify underlying action concepts.
@@
Predicting human motion in future frames is another widely used approach to interpret dynamic sequences and their corresponding actions. This technique allows us to gain valuable insights into the model’s training focus and potential biases towards the specific data it has been trained on.  [192] have successfully achieved an adaptive and concise representation for diverse graph structures of human motion when modeling the spatial-temporal actions using complex structured graphs. This graph-based approach incorporates the spatial and temporal dependencies by considering the sequential nature of the motion data which allows the model to effectively capture the temporal relationships between different frames in the motion sequence. Additionally, this framework can establish temporal dependencies in the subsequent frames, which can contribute to model interpretability.
A natural question that arises is if the decision-making process in deep learning models for action recognition in semiology can be interpretable. Considering the spread of deep processing for various medical applications, explainability and quantitative evaluation with a focus on usability by clinicians is crucial. A next step is to combine seizure video automated analyses with other investigations including anatomo-electroclinical correlations from stereoelectroencephalography (SEEG) data, which could bring new insights via multi-modal artificial intelligence approaches. Once sufficiently large and diverse human datasets are compiled for testing, such analyses could ultimately refine the power of automatically detected seizure video patterns to correctly classify seizure type and predict likely cerebral localisation.
@@ 6 Conclusion
Recent advances leveraging deep learning models for the detection and classification of seizure semiology show great promise. This survey provides insights from existing works and outlines several challenges and opportunities that will benefit the field and other medical applications. The proposed pipeline introduces a comprehensive integrated system, aiming to quantify and appropriately evaluate video recordings of seizures. As the field progresses, the exploration of multi-sensory methodologies capable of learning across semiology, electrical patterns, and neuroimaging for precise localization emerges as a key challenge for the neuro-engineering scientific community. While individual clinical expertise remains irreplaceable, the provision of more quantitative evidence holds the potential to enhance medical decision-making and ultimately elevate the standard of care for patients.

### 04 Mar 2024 v2

@@ 6.5 Explainable patient activity recognition based on interpretable models
One of the major hindrances in generating spatio-temporal interpretation outputs is the lack of a model-agnostic interpretation pipeline. This is challenging, as the spatio-temporal interpretation illustrates the significance of detecting biases in video data and action associations through deeper network analysis. The motion-guided sampler (MGSampler) [268] attempted to interpret the dynamic motion characteristics by leveraging motion as a fundamental signal for frame selection. It utilizes a frame selection approach that ensures an even distribution of frames across important segments, characterized by high motion salience, achieved through the accumulation of motion data. M-MiT [163] provides a more comprehensive method that is focused on interpreting action recognition models. This method is focused on examining these interpretable features and how they contribute to a network’s output. It reveals insights that might otherwise remain concealed without the capacity to identify underlying action concepts.
@@
Predicting human motion in future frames is another widely used approach to interpret dynamic sequences and their corresponding actions. This technique allows us to gain valuable insights into the model’s training focus and potential biases towards the specific data it has been trained on.  [214] have successfully achieved an adaptive and concise representation for diverse graph structures of human motion when modeling the spatio-temporal actions using complex structured graphs. This graph-based approach incorporates the spatial and temporal dependencies by considering the sequential nature of the motion data which allows the model to effectively capture the temporal relationships between different frames in the motion sequence. Additionally, this framework can establish temporal dependencies in the subsequent frames, which can contribute to model interpretability.
A natural question that arises is if the decision-making process in deep learning models for action recognition in semiology can be interpretable. Considering the spread of deep processing for various medical applications, explainability and quantitative evaluation with a focus on usability by clinicians is crucial [134]. A next step is to combine seizure video automated analyses with other investigations including anatomo-electroclinical correlations from stereoelectroencephalography (SEEG) data [22], which could bring new insights via multi-modal artificial intelligence approaches [123]. Once sufficiently large and diverse human datasets are compiled for testing, such analyses could ultimately refine the power of automatically detected seizure video patterns to correctly classify seizure type and predict likely cerebral localisation.
@@ 7 Conclusion
Recent advances leveraging deep learning models for the detection and classification of seizure semiology show great promise. This survey provides insights from existing works and outlines several challenges and opportunities that will benefit the field and other medical applications. The proposed pipeline introduces a comprehensive integrated system, aiming to quantify and appropriately evaluate video recordings of seizures. As the field progresses, the exploration of multi-sensory methodologies capable of learning across semiology, electrical patterns, and neuroimaging for precise localization emerges as a key challenge for the neuro-engineering scientific community. While clinical expertise at the individual and team levels remains irreplaceable, the provision of more quantitative, complementary evidence holds the potential to enhance medical decision-making and ultimately elevate the standard of care for patients.